Cove’s Cakes is a local bakery. Price and cost information follows:

Price per cake $ 13.81
Variable cost per cake
Ingredients 2.21
Direct labor 1.19
Overhead (box, etc.) 0.27
Fixed cost per month $ 4,765.80


Required:
1.
Determine Cove’s break-even point in units and sales dollars. (Round your Break-Even Units answer to the nearest whole number. Round your other intermediate calculations and sales dollars answer to 2 decimal places.)

   

2. Determine the bakery’s margin of safety if it currently sells 550 cakes per month. (Round your intermediate calculations to 2 decimals. Round the break-even units and final answer to nearest whole dollar.)

   

3. Determine the number of cakes that Cove must sell to generate $1,500 in profit. (Round your intermediate calculations to 2 decimal places and final answer to nearest whole number.)
Answer #1

Total variable cost=Ingredients+Direct labor+Overhead

=(2.21+1.19+0.27)=$3.67 per cake

Contribution margin=Sales-variable cost

=(13.81-3.67)=$10.14 per cake

1.Breakeven point=Fixed cost/Contribution margin

=(4765.8/10.14)=470 cakes

=(470*13.81)=$6490.70

2.Margin of safety=Total sales-Breakeven sales

=(550-470)=80 cakes

=(80*13.81)=$1104.80

=$1105(Approx).

3.Contribution margin required=Fixed cost+Target profits

=(4765.8+1500)=$6265.8

Hence cakes to be sold=$6265.8/10.14

=618 cakes(Approx).